OpenStudy (anonymous):

Check my Algebra answers please, :) Simplify the expression: 16 ÷ 8 • 2. A.4 B.32 C.1 D.16 [I think A] Simplify the expression: [(18-6)x3+1] x 7 A.43 B.259 C.7 D.336 [I think B] Evaluate the expression if a = 1 and b = 4. \[\frac{ 2(a+6b) }{ ab }\] A.12.5 B.14 C.6.5 D.8 [I think A] Translate the word of phrase into a variable expression. The product of a number and 5 is decreased by 2. A.\[\frac{ n }{ 5}-2\] B.\[2-\frac{ n }{ 5 }\] C. 2 – 5n D. 5n – 2 [I think its A] (ran out of room, The last question is below;

OpenStudy (anonymous):

Write a variable expression that represents the word problem. Gisele gave d dimes and n nickles to her brother. Which expression represents the amount in dollars that Gisele gave to her brother? A. (0.10 + 0.05) • nd B. 0.10d + 0.05n C. d • n D. d + n [I think its B]

zepdrix (zepdrix):

Your first and third ones look correct. The second question doesn't even have the correct responses, lemme show you:\[\Large\rm [(18-6)-3+1] \times 7\]\[\Large\rm [(12)-3+1] \times 7\]\[\Large\rm [9+1] \times 7\]\[\Large\rm [10] \times 7\]\[\Large\rm 70\]

zepdrix (zepdrix):

The product of a number and 5 is decreased by 2. So you've got the subtraction correct, you're subtracting 2 from something. But you're subtracting 2 from a `product`, not from a quotient`. So 5 should be multiplying n.

zepdrix (zepdrix):

5th question looks correct.

OpenStudy (anonymous):

Wait I have more :( sorry Which expression completes this equation? 10 • 2 + 15 ÷ 3 = ? A. 10 • 17 ÷ 3 B. 35 – [6 + (3 – 1)] C. 5 • [(7 – 3) + 1] D. 9 • 2 + 1 – 3 [I think C] Which value of x makes the open sentence true? 7 – x = 3x + 5 A. x = 1.5 B. x = 3 C. x = 0.5 D. x = 2 [I think C] Which equation represents this situation? Charlie worked 1.5 hours less than Dirk. Together, they worked a total of 25 hours. Dirk worked h hours. A. h – 1.5 = 25 B. h + (h + 1.5) = 25 C. h – 2.5 = 25 D. h + (h – 1.5) = 25 [I think D] Nia wants to place grouping symbols in the expression 35 ÷ 7 – 5 + 4 so that it has a value of 21.5. Which shows the correct placing of the grouping symbols? A. (35 ÷ 7) – 5 + 4 B. 35 ÷ 7 – (5 + 4) C. 35 ÷ (7 – 5) + 4 D. 35 ÷ (7 – 5 + 4) [I think C] Thank you so much for the help =)

zepdrix (zepdrix):

10 • 2 + 15 ÷ 3 = 25 Ok good job, option C seems to be giving 25 as well. 7 – x = 3x + 5 x=1/2, good. `h` + `(h – 1.5)` = `25` `Dirk's hours` + `Charlie's Hours` = `Total Hours`. Ok good on this one also. C. 35 ÷ (7 – 5) + 4 Yes good job! :)

OpenStudy (anonymous):

I messed up on question 2 :( , I accidentally made it "-3" instead of "x3" sorry

zepdrix (zepdrix):

\[\Large\rm [(18-6)\times3+1] \times 7\]\[\Large\rm [(12)\times3+1] \times 7\]\[\Large\rm [36+1] \times 7\]\[\Large\rm [37] \times 7\]\[\Large\rm 259\]Ok great! Your answer looks there.

OpenStudy (anonymous):

yeah, thanks you soo much =D

zepdrix (zepdrix):

c: